# On the Role of Entity and Event Level Conceptualization in Generalizable Reasoning: A Survey of Tasks, Methods, Applications, and Future Directions

### v1

@@ 6.3 Conclusions
The main limitations of our survey are two-fold. First, due to the vast amount of literature on conceptualization and conceptual knowledge across various datasets, we only cover the most representative works that stand out for their exceptional value and uniqueness in our taxonomy. Most of the papers are sourced from ACL Anthology [^1] https://aclanthology.org/, ACM Digital Library [^2] https://dl.acm.org/, and proceedings of leading artificial intelligence and machine learning conferences. Consequently, it is possible that some important works were not included. Second, our survey specifically focuses on entity and event level conceptualization, leaving document/paragraph level and system level conceptualization unaddressed. However, it is impossible to survey everything within one single submission. Future research can expand the scope of our survey to include more types of conceptualizations and modalities, such as categorization in the vision modality Chen and Wang (2004).
@@ A.1 Extraction Based Methods
While these methods have been successful in extracting concepts and relationships from text, they have several limitations. First, they are heavily dependent on the quality of the text and the predefined list of concepts. If the text is noisy or contains many irrelevant words, the performance of these methods can degrade significantly, and the resulting extracted concepts may also tend to be noisy. Second, it’s important to note that these methods primarily rely on parsing or pattern matching techniques on text and do not capture semantic information from the text. This potentially makes extracted concepts represented as isolated entities without any context or relationships and could result in mis-extraction of concepts or relationships, especially when the text contains ambiguous or polysemous words. For example, the word “bank” can refer to a financial institution, a river bank, or a memory bank, and without proper context, it’s difficult to determine the correct meaning of it, thus leading to incorrect concept extraction. A low-performance parser, if wrongly parsing these words, may also lead to noisy results. Lastly, these methods are not able to generalize well to unseen concepts or text patterns that are not present in the predefined list of concepts. This limits their applicability to new domains or tasks that require the extraction of novel concepts or relationships. For example, to extract concepts from medical or legal domain text, specific patterns or extraction rules need to be designed, which may not be present when extracting normal conversational text.
@@ A.3.2 Zero-Shot Generative Methods
Zero-shot generative methods aim to generate the desired output for any task’s input without any task-specific fine-tuning. A very representative example of such generative models is the recently popularized LLMs OpenAI (2022, 2023); Touvron et al. (2023a, b); Mesnard et al. (2024); Reid et al. (2024). These models have been pre-trained on very large corpora, including those from the web, Wikipedia, books, and more, and have shown strong performance in various natural language processing tasks, including text generation Maynez et al. (2023); Chen et al. (2024), temporal reasoning Tan et al. (2023); Yuan et al. (2024), causal reasoning Chan et al. (2024a); Dalal et al. (2023); Jin et al. (2023), commonsense reasoning Jain et al. (2023); Bian et al. (2023); Fang et al. (2021b, a); Deng et al. (2023), logical reasoning Wang et al. (2023d, e, 2021); Bai et al. (2023), and more Qin et al. (2023); Cheng et al. (2023); Chan et al. (2024b).

### v2

@@ 6.3 Conclusions
The main limitations of our survey are two-fold. First, due to the vast amount of literature on conceptualization and conceptual knowledge across various datasets, we only cover the most representative works that stand out for their exceptional value and uniqueness in our taxonomy. Most of the papers are sourced from ACL Anthology [^1] https://aclanthology.org/, ACM Digital Library [^2] https://dl.acm.org/, and proceedings of leading artificial intelligence and machine learning conferences. Consequently, it is possible that some other related works are not included, but we aim to cover them in future versions. Second, our survey specifically focuses on entity and event level conceptualization, leaving document/paragraph level and system level conceptualization unaddressed. However, it is impossible to survey everything within one single submission. Future research can expand the scope of our survey to include more types of conceptualizations and modalities, such as categorization in the vision modality Chen and Wang (2004).
@@
Acknowledgments
We thank the anonymous reviewers and the area chair for their constructive comments. The authors of this paper were supported by the ITSP Platform Research Project (ITS/189/23FP) from the ITC of Hong Kong, SAR, China, as well as the AoE (AoE/E-601/24-N), the RIF (R6021-20), and the GRF (16205322) from the RGC of Hong Kong, SAR, China.
@@ A.1 Extraction Based Methods
While these methods have been successful in extracting concepts and relationships from text, they have several limitations. First, they are heavily dependent on the quality of the text and the predefined list of concepts. If the text is noisy or contains many irrelevant words, the performance of these methods can degrade significantly, and the resulting extracted concepts may also tend to be noisy. Second, it’s important to note that these methods primarily rely on parsing or pattern matching techniques on text and do not capture semantic information from the text Li et al. (2025). This potentially makes extracted concepts represented as isolated entities without any context or relationships and could result in mis-extraction of concepts or relationships, especially when the text contains ambiguous or polysemous words. For example, the word “bank” can refer to a financial institution, a river bank, or a memory bank, and without proper context, it’s difficult to determine the correct meaning of it, thus leading to incorrect concept extraction. A low-performance parser, if wrongly parsing these words, may also lead to noisy results. Lastly, these methods are not able to generalize well to unseen concepts or text patterns that are not present in the predefined list of concepts. This limits their applicability to new domains or tasks that require the extraction of novel concepts or relationships. For example, to extract concepts from medical or legal domain text, specific patterns or extraction rules need to be designed, which may not be present when extracting normal conversational text.
@@ A.3.2 Zero-Shot Generative Methods
Zero-shot generative methods aim to generate the desired output for any task’s input without any task-specific fine-tuning. A very representative example of such generative models is the recently popularized LLMs OpenAI (2022, 2023); Touvron et al. (2023a, b); Mesnard et al. (2024); Reid et al. (2024). These models have been pre-trained on very large corpora, including those from the web, Wikipedia, books, and more, and have shown strong performance in various natural language processing tasks, including text generation Maynez et al. (2023); Chen et al. (2024); Wang et al. (2025a), temporal reasoning Tan et al. (2023); Yuan et al. (2024), causal reasoning Chan et al. (2024a); Dalal et al. (2023); Jin et al. (2023), commonsense reasoning Jain et al. (2023); Bian et al. (2023); Fang et al. (2021b, a); Deng et al. (2023), logical reasoning Wang et al. (2023d, e, 2021); Bai et al. (2023), and more Qin et al. (2023); Cheng et al. (2023); Chan et al. (2024b); Wang et al. (2025b).